OpenStudy (anonymous):

find the derivative of the function…. 3(4-x^2)^5 can someone please explain how to do this step by step…I need at least an example so I can do my other problems that have to do with finding derivatives….PLEASE…And NO I DONT JUST WANT THE ANSWER. I want to be able to understand this…. Please help...

OpenStudy (anonymous):

d (x^n)/ dx= ?

OpenStudy (anonymous):

u know the derivative of x^n w.r.t. x?

OpenStudy (anonymous):

nooooo?

OpenStudy (anonymous):

derivative of x^n w.r.t. x = d (x^n)/ dx = n* x^(n-1) fine?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u have to find the derivative of 3(4-x^2)^5 wrt x u know chain rule?

OpenStudy (anonymous):

if no...forget it for now.

OpenStudy (anonymous):

yes.. i do kinda

OpenStudy (anonymous):

derivative of 3(4-x^2)^5 wrt x suppose (4-x^2)=t so now we have to find the derivative of 3 t^5 wrt x. d(3 t^5)/ dx= 3 d(t^5)/dx, 3 is a constant ...so pulled it outside.

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

because derivative of a constant term is 0 so 3 is pulled out outside.

OpenStudy (anonymous):

does it make sense?

mathslover (mathslover):

jessiedeee , Did you get it ?

OpenStudy (anonymous):

sorry, I had to go…. but @akash123 yes It makes sense...

mathslover (mathslover):

Good!